If f(x+y)=f(x)+f(y) and f(1)=17, find f(2), f(3), and f(4). Is f(x+y)=f(x)+f(y) for all functions?
f(2)=
f(3)=
f(4)=

Answer by ikleyn(53763)

.

then  f(2) = f(1 + 1) = f(1) + f(1) = 17 + 17 = 34.     ANSWER


      f(3) = f(1 + 2) = f(1) + f(2) = 17 + 34 = 51.     ANSWER


      f(4) = f(2 + 2) = f(2) + f(2) = and you continue from this point.

------------

For the last question, the answer is "NO".

As an counterexample, consider f(x) = x^2.


/\/\/\/\/\/\/

I ask the visitors to post their "THANKS" for my responses,

in order for I see that I work for civilized humans . . .


. . . and not for idiots . . .


Otherwise, I will write to your MANAGERS asking them to check you for your professional qualification
to occupy your position in this project . . .